# EHRAgent: Code Empowers Large Language Models for Complex Tabular Reasoning on Electronic Health Records

### v1

@@ Acknowledgments
This research has been partially supported by Accelerate Foundation Models Academic Research Initiative from Microsoft Research, Wallace H. Coulter Distinguished Faculty Fellowship, Petit Institute Faculty Fellowship, and research funding from Amazon and Microsoft Research to Professor May D. Wang.
Appendix A Dataset Details
@@ A.1 Task Details
We evaluate EHRAgent on a text-to-SQL medical question answering (QA) benchmark (Lee et al., 2022), EHRSQL [^3] https://github.com/glee4810/EHRSQL, built upon structured EHRs from MIMIC-III and eICU. EHRSQL serves as a text-to-SQL benchmark for assessing the performance of medical QA models, specifically focusing on generating SQL queries for addressing a wide range of real-world questions gathered from over 200 hospital staff. Questions within EHRSQL, ranging from simple data retrieval to complex operations such as calculations, reflect the diverse and complex clinical tasks encountered by front-line healthcare professionals.
@@ A.2 Question Complexity Level
We categorize input queries from EHRSQL (Lee et al., 2022) into various complexity levels (levels I-IV for MIMIC-III and levels I-III for eICU) based on the number of tables involved in solution generation. For example, given the question ‘how many patients were given temporary tracheostomy?’, the complexity level is categorized as II, indicating that we need to extract information from two tables (admission and procedure) to generate the solution. Furthermore, we also conduct a performance analysis (see Figure 3) based on additional evaluation metrics related to question complexity, including (1) number of elements (i.e., slots) in each question and (2) the number of columns involved in each solution. Specifically, elements refer to the slots within each template that can be populated with pre-defined values or database records.
@@ A.3 MIMIC-III
Similar to MIMIC-III, eICU (Pollard et al., 2018) [^5] https://physionet.org/content/eicu-crd/2.0/ includes over 200,000 admissions from multiple critical care units across the United States in 2014 and 2015. It contains deidentified administrative information following US Health Insurance Portability and Accountability Act (HIPAA) standard, and structured clinical data, including vital signs, laboratory measurements, medications, treatment plan, admission diagnosis, and medical history.
Appendix B Additional Implementation Details
B.1 Tool Sets Details
@@
B.2 Comparison of Baselines
We compare baselines and EHRAgent on the inclusion of different components in Table 4.
Baselines Tool Use Code Interface Environment Feedback Debugging Error Exploration Medical Knowledge Long-term Memory
@@ B.2 Comparison of Baselines
AutoGen (Wu et al., 2023b)
@@
B.3 Hardware and Software Details
@@
All experiments are conducted on CPU: Intel(R) Core(TM) i7-5930K CPU @ 3.50GHz and GPU: NVIDIA GeForce RTX A5000 GPUs, using Python 3.9 and AutoGen 0.2.0777https://github.com/microsoft/autogen.
B.4 Code Generation Details
@@
Given that the majority of LLMs have been pre-trained on Python code snippets (Gao et al., 2023), and Python’s inherent modularity aligns well with tool functions, we choose Python 3.9 as the primary language for interaction coding between the LLM agent and the code executor.
B.5 Prompt Details
@@
In the subsequent subsections, we detail the prompt templates employed in EHRAgent. The complete version of prompts is available at our GitHub repository due to space limitations.
@@
Appendix C Additional Experimental Results
C.1 Additional Ablation Studies
@@
$\diamond$ Medical Knowledge Integration. Out of all the components, the medical knowledge injection module mainly exhibits its benefits in challenging tasks. These tasks often involve more tables and require a deeper understanding of domain knowledge to associate items with their corresponding tables.
$\diamond$ Interactive Coding. The interactive coding interface is the most significant contributor to the performance gain across all complexity levels. This verifies the importance of utilizing the code interface for planning instead of natural languages, which enables the model to avoid overly complex contexts and thus leads to a substantial increase in the completion rate. Additionally, the code interface also allows the debugging module to refine the planning with execution feedback, improving the efficacy of the planning process.
$\diamond$ Debugging Module. The ‘rubber duck’ debugging module enhances the performance by guiding the LLM agent to figure out the underlying reasons for the error messages. This enables EHRAgent to address the intrinsic error that occurs in the original reasoning steps.
$\diamond$ Long-term Memory. Following the reinforcement learning setting (Sun et al., 2023; Shinn et al., 2023), the long-term memory mechanism improves performance by justifying the necessity of selecting the most relevant demonstrations for planning. In order to simulate the scenario where the ground truth annotations (i.e., rewards) are unavailable, we further evaluate the effectiveness of the long-term memory on the completed cases in Table 6, regardless of whether they are successful or not. The results indicate that the inclusion of long-term memory with completed cases increases the completion rate but tends to reduce the success rate across most difficulty levels, as some incorrect cases might be included as the few-shot demonstrations. Nonetheless, it still outperforms the performance without long-term memory, confirming the effectiveness of memory mechanism.
@@
 w/o medical knowledge 36.75 28.39 6.25 30.17 47.24
@@
 w/o debugging 50.60 46.98 12.50 47.07 70.86
 w/o long-term memory 52.41 44.22 18.75 45.69 78.97
@@
Appendix D Additional Empirical Analysis
D.1 Additional Question Complexity Analysis
We further analyze the model performance by considering various measures of question complexity based on the number of elements in questions, and the number of columns involved in solutions, as shown in Figure 3. Incorporating more elements requires the model to either perform calculations or utilize domain knowledge to establish connections between elements and specific columns. Similarly, involving more columns also presents a challenge for the model in accurately locating and associating the relevant columns. We notice that both EHRAgent and baselines generally exhibit lower performance on more challenging tasks [^8] Exceptions may exist when considering questions of seven elements in Figures 3(a) and 3(b), as it comprises only eight samples and may not be as representative.. Notably, our model consistently outperforms all the baseline models across all levels of difficulty. Specifically, for those questions with more than 10 columns, the completion rate of those open-loop baselines is very low (less than 20%), whereas EHRAgent can still correctly answer around 50% of queries, indicating the robustness of EHRAgent in handling complex queries with multiple elements.
D.2 Additional Error Analysis
We conducted a manual examination to analyze all incorrect cases generated by EHRAgent in MIMIC-III. Figure 6 illustrates the percentage of each type of errors frequently encountered during solution generation:

### v3

@@ Acknowledgments
We thank the anonymous reviewers and area chairs for their valuable feedback. This research was partially supported by Accelerate Foundation Models Academic Research Initiative from Microsoft Research. This research was also partially supported by the National Science Foundation under Award Number 2319449 and Award Number 2312502, the National Institute Of Diabetes And Digestive And Kidney Diseases of the National Institutes of Health under Award Number K25DK135913, the Emory Global Diabetes Center of the Woodruff Sciences Center, Emory University.
Appendix A Dataset and Task Details
@@ A.1 Task Details
We evaluate EHRAgent on three publicly available EHR datasets from two text-to-SQL medical question answering (QA) benchmarks (Lee et al., 2022), EHRSQL [^7] https://github.com/glee4810/EHRSQL and TREQS [^8] https://github.com/wangpinggl/TREQS, built upon structured EHRs from MIMIC-III and eICU. EHRSQL and TREQS serve as text-to-SQL benchmarks for assessing the performance of medical QA models, specifically focusing on generating SQL queries for addressing a wide range of real-world questions gathered from over 200 hospital staff. Questions within EHRSQL and TREQS, ranging from simple data retrieval to complex operations such as calculations, reflect the diverse and complex clinical tasks encountered by front-line healthcare professionals. Dataset statistics are available in Table 3.
Dataset # Examples # Table # Row/Table # Table/Q
MIMIC-III 580 17 81k 2.52
eICU 580 10 152k 1.74
TREQS 996 5 498k 1.48
Average 718.7 10.7 243.7k 1.91
Table 3: Dataset statistics.
@@ A.2 Question Complexity Level
We categorize input queries into various complexity levels (levels I-IV for MIMIC-III and levels I-III for eICU and TREQS) based on the number of tables involved in solution generation. For example, given the question ‘How many patients were given temporary tracheostomy?’, the complexity level is categorized as II, indicating that we need to extract information from two tables (admission and procedure) to generate the solution. Furthermore, we also conduct a performance analysis (see Figure 4) based on additional evaluation metrics related to question complexity, including (1) the number of elements (i.e., slots) in each question and (2) the number of columns involved in each solution. Specifically, elements refer to the slots within each template that can be populated with pre-defined values or database records.
@@ A.3 MIMIC-III
Similar to MIMIC-III, eICU (Pollard et al., 2018) [^10] https://physionet.org/content/eicu-crd/2.0/ includes over 200,000 admissions from multiple critical care units across the United States in 2014 and 2015. It contains deidentified administrative information following the US Health Insurance Portability and Accountability Act (HIPAA) standard and structured clinical data, including vital signs, laboratory measurements, medications, treatment plans, admission diagnoses, and medical histories.
A.5 TREQS
TREQS (Wang et al., 2020) is a healthcare question and answering benchmark that is built upon the MIMIC-III (Johnson et al., 2016) dataset. In TREQS, questions are generated automatically using pre-defined templates with the text-to-SQL task. Compared to the MIMIC-III dataset within the EHRSQL (Lee et al., 2022) benchmark, TREQS has a narrower focus in terms of the types of questions and the complexity of SQL queries. Specifically, it is restricted to only five tables but includes a significantly larger number of records (Table 3) within each table.
Appendix B Tool Set Details
@@
Appendix C Baseline Details
All the methods, including baselines and EHRAgent, share the same (1) tool definitions, (2) table meta information, and (3) few-shot demonstrations in the prompts by default. The only difference is the prompting style or technical differences between different methods, which guarantees a fair comparison among all baselines and EHRAgent. Table 4 summarizes the inclusion of different components in both baselines and ours.
Baselines Tool Use Code Interface Environment Feedback Debugging Error Exploration Medical Information Long-term Memory
   w/o Code Interface
@@ Appendix C Baseline Details
DIN-SQL (Pourreza and Rafiei, 2023)
Self-Debugging (Chen et al., 2024)
AutoGen (Wu et al., 2023)
@@
$\bullet$ Baselines w/o Code Interface. LLMs without a code interface rely purely on natural language-based planning capabilities.
$\diamond$ CoT (Wei et al., 2022): CoT enhances the complex reasoning capabilities of original LLMs by generating a series of intermediate reasoning steps.
$\diamond$ Self-Consistency (Wang et al., 2023d): Self-consistency improves CoT by sampling diverse reasoning paths to replace the native greedy decoding and select the most consistent answer.
$\diamond$ Chameleon (Lu et al., 2023): Chameleon employs LLMs as controllers and integrates a set of plug-and-play modules, enabling enhanced reasoning and problem-solving across diverse tasks.
$\diamond$ ReAct (Yao et al., 2023b): ReAct integrates reasoning with tool use by guiding LLMs to generate intermediate verbal reasoning traces and tool commands.
$\diamond$ Reflexion (Shinn et al., 2023): Reflexion leverages verbal reinforcement to teach LLM-based agents to learn from linguistic feedback from past mistakes.
$\bullet$ Baselines w/ Code Interface. LLMs with a code interface enhance the inherent capabilities of LLMs by enabling their interaction with programming languages and the execution of code. In accordance with their default configuration, we present a summary of the utilization of programming languages in various baselines in Table 5. Additionally, we provide a detailed explanation of the programming language selection in EHRAgent in Appendix D.
$\diamond$ LLM2SQL (Nan et al., 2023): LLM2SQL augments LLMs with a code interface to generate SQL queries for retrieving information from EHRs for question answering.
$\diamond$ DIN-SQL (Pourreza and Rafiei, 2023): Compared to LLM2SQL, DIN-SQL further breaks down a complex problem into several sub-problems and feeding the solutions of those sub-problems into LLMs, effectively improving problem-solving performance.
$\diamond$ Self-Debugging (Chen et al., 2024): Self-Debugging teaches LLMs to debug by investigating execution results and explaining the generated code in natural language.
$\diamond$ AutoGen (Wu et al., 2023): AutoGen unifies LLM-based agent workflows as multi-agent conversations and uses the code interface to encode interactions between agents and environments. We follow the official tutorial [^12] https://microsoft.github.io/autogen/docs/Use-Cases/agent_chat/ for the implementation of AutoGen. Specifically, we utilize the built-in AssistantAgent and UserProxyAgent within AutoGen to serve as the LLM agent and the code executor, respectively. The AssistantAgent is configured in accordance with AutoGen’s tailored system prompts, which are designed to direct the LLM to (1) propose code within a coding block as required, (2) refine the proposed code based on execution outcomes, and (3) append a specific code, "TERMINATE", to conclude the response for terminating the dialogue. The UserProxyAgent functions as a surrogate for the user, extracting and executing code from the LLM’s responses in a local environment. Subsequently, it relays the execution results back to the LLM. In instances where code is not detected, a standard message is dispatched instead. This arrangement facilitates an automated dialogue process, obviating the need for manual tasks such as code copying, pasting, and execution by the user, who only needs to initiate the conversation with an original query.
Baselines # Language
LLM2SQL (Nan et al., 2023) SQL
DIN-SQL (Pourreza and Rafiei, 2023) SQL
Self-Debugging (Chen et al., 2024) SQL
AutoGen (Wu et al., 2023) Python
EHRAgent (Ours) Python
Table 5: Comparison of baselines and EHRAgent on the selection of primary programming languages.
Appendix D Selection of Primary Programming Language
In our main experiments, we concentrate on three SQL-based EHR QA datasets to assess EHRAgent in comparison with other baselines. Nevertheless, we have opted for Python as the primary programming language for EHRAgent, rather than SQL [^13] We include an empirical analysis in Appendix G.3 to further justify the selection of Python as the primary programming language for EHRAgent.. The primary reasons for choosing Python instead of SQL to address medical inquiries based on EHRs are outlined below:
Python Enables the External Tool-Use.
Using alternative programming languages, such as SQL, can result in LLM-based agents becoming unavailable to external tools or functions. The primary contribution of EHRAgent is to develop a code-empowered agent capable of generating and executing code-based plans to solve complex real-world clinic tasks. In general, the SQL language itself is incapable of calling API functions. For example, EHRXQA (Bae et al., 2023) can be considered as an LLM agent that generates a solution plan in NeuralSQL (not SQL). This agent is equipped with two tools: a pre-trained Visual Question Answering (VQA) model called FUNC_VQA, and a SQL interpreter. Similar to EHRAgent, it also relies on a non-SQL language and includes an SQL interpreter as a tool. Compared with NeuralSQL in EHRXQA (Bae et al., 2023), Python in EHRAgent can be directly executed, while NeuralSQL requires additional parsing.
Python Enables the Integration of SQL Tool Function.
Python provides excellent interoperability with various databases and data formats. It supports a wide range of database connectors, including popular relational databases such as PostgreSQL, MySQL, and SQLite, as well as non-relational databases like MongoDB. This interoperability ensures that EHRAgent can seamlessly interact with different EHR systems and databases. Although our proposed method primarily relies on generating and executing Python code, we do not prohibit EHRAgent from utilizing SQL to solve problems. In our prompts and instructions, we also provide the ’SQLInterpreter’ tool function for the agent to perform relational database operations using SQL. Through our experiments, we have observed that EHRAgent is capable of combining results from Python code and SQL commands effectively. For instance, when presented with the question, “Show me patient 28020’s length of stay of the last hospital stay.”, EHRAgent will first generate SQL command admit_disch_tuple = SQLInterpreter (ŚELECT ADMITTIME, DISCHTIME FROM admissions WHERE SUBJECT_ID=28020 ORDER BY ADMITTIME DESC LIMIT 1)́ and execute it to obtain the tuples containing the patient’s admission and discharge times. It will then employ Python code along with the built-in date-time function to calculate the duration of the last stay tuple.
Python Enables a More Generalizable Framework.
EHRAgent is a generalizable LLM agent empowered with a code interface to autonomously generate and execute code as solutions to given problems. While Section 4 focuses on the challenging multi-tabular reasoning task within EHRs for evaluation, the Python-based approach has the potential to be generalized to other tasks (e.g., risk prediction tasks based on EHRs) or even multi-modal clinical data and be integrated with additional tool-sets in the future. In contrast, other languages like SQL are limited to database-related operations.
Python is More Flexible in Extension.
Python is a general-purpose programming language that offers greater flexibility compared to SQL. It enables the implementation of complex logic and algorithms, which may be necessary for solving certain types of medical questions that require more than simple database queries. Python is also a highly flexible programming language that offers extensive capabilities through its libraries and frameworks, making it suitable for handling a wide range of programming tasks, including database operations. In contrast, SQL is only applicable within relational databases and does not provide the same level of flexibility and extension. This attribute is particularly important to LLM-based agents, as they can leverage both existing Python libraries and custom-defined functions as tools to solve complex problems that are inaccessible for and beyond the scope of SQL.
Python Includes More Extensive Resources for Pre-Training.
Python has a large and active community of developers and researchers. This community contributes to the development of powerful libraries, frameworks, and tools that can be leveraged in EHRAgent. The extensive documentation, tutorials, and forums available for Python also provide valuable resources for troubleshooting and optimization. Github repositories are one of the most extensive sources of code data for state-of-the-art language models (i.e., LLMs), such as GPTs. Python is the most widely used coding language on Github [^14] https://madnight.github.io/githut/#/pull_requests/2023/1. In addition, Python is known for its readability and maintainability. The clean and expressive syntax of Python makes it easier for researchers and developers to understand, modify, and extend the codebase of EHRAgent. This is particularly important when extended to real-world clinical research and practice, where the system may need to be updated frequently to incorporate new knowledge and adapt to evolving requirements.
Appendix E Additional Implementation Details
E.1 Hardware and Software Details
@@
All experiments are conducted on CPU: Intel(R) Core(TM) i7-5930K CPU @ 3.50GHz and GPU: NVIDIA GeForce RTX A5000 GPUs, using Python 3.9 and AutoGen 0.2.0151515https://github.com/microsoft/autogen.
E.2 Data Preprocessing Details
During the data pre-processing stage, we create EHR question-answering pairs by considering text queries as questions and executing SQL commands in the database to automatically generate the corresponding ground-truth answers. We filter out samples containing unexecutable SQL commands or yielding empty results throughout this process.
E.3 Code Generation Details
@@
Given that the majority of LLMs have been pre-trained on Python code snippets (Gao et al., 2023), and Python’s inherent modularity aligns well with tool functions, we choose Python 3.9 as the primary coding language for interaction coding and AutoGen 0.2.0 Wu et al. (2023) as the interface for communication between the LLM agent and the code executor.
E.4 Selection of Initial Set of Demonstrations
The initial set of examples is collected manually, following four criteria: (1) using the same demonstrations across all the baselines; (2) utilizing all the designed tools; (3) covering as many distinct tables as possible; and (4) including examples in different styles of questions. With these criteria in mind, we manually crafted four demonstrations for each dataset. To ensure a fair comparison, we use the same initial four-shot demonstrations ($K=4$) for all baselines and EHRAgent, considering the maximum length limitations of input context in baselines like ReAct (Yao et al., 2023b) and Chameleon (Lu et al., 2023).
E.5 Evaluation Metric Details
Our main evaluation metric is the success rate (SR), quantifying the percentage of queries that the model successfully handles. In addition, we leverage completion rate (CR) as a side evaluation metric to represent the percentage of queries for which the model is able to generate executable plans, regardless of whether the results are correct. Specifically, following existing LLM-based agent studies (Xu et al., 2023; Kirk et al., 2024), we use CR to assess the effectiveness of LLM-based agents in generating complete executable plans without execution errors. One of our key components in EHRAgent is interactive coding with environmental feedback. By using CR, we can demonstrate that our proposed EHRAgent, along with other baselines that incorporate environmental feedback (e.g., ReAct (Yao et al., 2023b), Reflexion (Shinn et al., 2023), Self-Debugging (Chen et al., 2024), and AutoGen (Wu et al., 2023)), has a stronger capability (higher CR) in generating complete executable plans without execution errors, compared to baselines without environmental feedback (e.g., CoT (Wei et al., 2022), Self-Consistency (Wang et al., 2023d), Chameleon (Lu et al., 2023), and LLM2SQL (Nan et al., 2023)).
E.6 EHR Metadata Details
$\diamond$ MIMIC-III.
 <MIMIC_III> Metadata
 Read the following data descriptions, generate
     the background knowledge as the context
     information that could be helpful for
     answering the question.
 (1) Tables are linked by identifiers which
     usually have the suffix ’ID’. For example,
     SUBJECT_ID refers to a unique patient,
     HADM_ID refers to a unique admission to
     the hospital, and ICUSTAY_ID refers to a
     unique admission to an intensive care unit.
 (2) Charted events such as notes, laboratory
     tests, and fluid balance are stored in a
     series of ’events’ tables. For example the
     outputevents table contains all measurements
     related to output for a given patient,
     while the labevents table contains
     laboratory test results for a patient.
 (3) Tables prefixed with ’d_’ are dictionary
     tables and provide definitions for
     identifiers.c For example, every row of
     chartevents is associated with a single
     ITEMID which represents the concept
     measured,  but it does not contain the
     actual name of the measurement. By joining
     chartevents and d_items on ITEMID, it is
     possible to identify the concept represented
     by a given ITEMID.
 (4) For the databases, four of them are used to
     define and track patient stays: admissions,
     patients,icustays, and transfers. Another
     four tables are dictionaries for cross-
     referencing codes against their respective
     definitions: d_icd_diagnoses,
     d_icd_procedures, d_items, and d_labitems.
     The remaining tables, including chartevents,
     cost, inputevents_cv, labevents,
     microbiologyevents, outputevents,
     prescriptions, procedures_icd, contain data
     associated with patient care, such as
     physiological measurements, caregiver
     observations, and billing information.
$\diamond$ eICU.
 <eICU> Metadata
 Read the following data descriptions, generate
     the background knowledge as the context
     information that could be helpful for
     answering the question.
 (1) Data include vital signs, laboratory
     measurements, medications, APACHE components,
     care plan information, admission diagnosis,
     patient history, time-stamped diagnoses
     from a structured problem list, and
     similarly chosen treatments.
 (2) Data from each patient is collected into a
     common warehouse only if certain interfaces
     are available. Each interface is used to
     transform and load a certain type of data:
     vital sign interfaces incorporate vital
     signs, laboratory interfaces provide
     measurements on blood samples, and so on.
 (3) It is important to be aware that different
     care units may have different interfaces in
     place, and that the lack of an interface
     will result in no data being available for
     a given patient, even if those measurements
     were made in reality. The data is provided
     as a relational database, comprising
     multiple tables joined by keys.
 (4) All the databases are used to record
     information associated to patient care,
     such as allergy, cost, diagnosis,
     intakeoutput, lab, medication, microlab,
     patient, treatment, vitalperiodic.
$\diamond$ TREQS.
 <TREQS> Metadata
 Read the following data descriptions, generate
     the background knowledge as the context
     information that could be helpful for
     answering the question.
 (1) The database contains five categories of
     information for patients, including
     demographics, laboratory tests, diagnosis,
     procedures and prescriptions, and prepared
     a specific table for each category
     separately.
 (2) These tables compose a relational patient
     database where tables are linked through
     patient ID and admission ID.
E.7 Prompt Details
@@
In the subsequent subsections, we detail the prompt templates employed in EHRAgent. The complete version of the prompts is available at our code repository due to space limitations.
@@
Appendix F Additional Experimental Results
F.1 Effect of Base LLMs
Table 6 presents a summary of the experimental results obtained from EHRAgent and all baselines using a different base LLM, GPT-3.5-turbo (0613). The results clearly demonstrate that EHRAgent continues to outperform all the baselines, achieving a performance gain of 6.72%. This highlights the ability of EHRAgent to generalize across different base LLMs as backbone models. When comparing the experiments conducted with GPT-4 (Table 1), the performance of both the baselines and EHRAgent decreases. This can primarily be attributed to the weaker capabilities of instruction-following and reasoning in GPT-3.5-turbo.
Dataset ($\rightarrow$) MIMIC-III
Complexity Level ($\rightarrow$) I II III IV All
Methods ($\downarrow$) /Metrics ($\rightarrow$) SR. SR. CR.
w/o Code Interface
CoT (Wei et al., 2022) 23.16 10.40 2.99 1.71 8.62 41.55
Self-Consistency (Wang et al., 2023d) 25.26 11.88 4.19 2.56 10.52 47.59
Chameleon (Lu et al., 2023) 27.37 11.88 3.59 2.56 11.21 47.59
ReAct (Yao et al., 2023b) 26.32 10.89 3.59 3.42 9.66 61.21
Reflexion (Shinn et al., 2023) 30.53 12.38 9.58 8.55 13.28 66.72
w/ Code Interface
LLM2SQL (Nan et al., 2023) 21.05 15.84 4.19 2.56 10.69 59.49
Self-Debugging (Chen et al., 2024) 36.84 33.66 22.75 16.24 27.59 72.93
AutoGen (Wu et al., 2023) 28.42 25.74 13.17 10.26 19.48 52.42
EHRAgent (Ours) 43.16 42.57 29.94 18.80 34.31 78.80
Table 6: Experimental results of success rate (i.e., SR.) and completion rate (i.e., CR.) on MIMIC-III using GPT-3.5-turbo as the base LLM. The complexity of questions increases from Level I (the simplest) to Level IV (the most difficult).
F.2 Additional Ablation Studies
@@
$\diamond$ Medical Information Integration. Out of all the components, the medical knowledge injection module mainly exhibits its benefits in challenging tasks. These tasks often involve more tables and require a deeper understanding of domain knowledge to associate items with their corresponding tables.
$\diamond$ Long-term Memory. Following the reinforcement learning setting (Sun et al., 2023; Shinn et al., 2023), the long-term memory mechanism improves performance by justifying the necessity of selecting the most relevant demonstrations for planning. In order to simulate the scenario where the ground truth annotations (i.e., rewards) are unavailable, we further evaluate the effectiveness of the long-term memory on the completed cases in Table 8, regardless of whether they are successful or not. The results indicate that the inclusion of long-term memory with completed cases increases the completion rate but tends to reduce the success rate across most difficulty levels, as some incorrect cases might be included as the few-shot demonstrations. We have also performed multi-round experiments with shuffled order and observed that the order had almost no influence on the final performance in all three datasets. Nonetheless, it still outperforms the performance without long-term memory, confirming the effectiveness of the memory mechanism.
$\diamond$ Interactive Coding. For the ablation study setting of EHRAgent w/o interactive coding, we directly chose CoT (Wei et al., 2022) as the backbone, where we deteriorate from generating code-based plans to natural language-based plans. Once the steps are generated, we execute them in a step-by-step manner and obtain error information from the tool functions. By combining the error messages with tool definitions and language-based plans, we are still able to prompt the LLMs to deduce the most probable underlying cause of the error. The medical information injection and long-term memory components remain unchanged from the original EHRAgent. From the ablation studies, we can observe that the interactive coding interface is the most significant contributor to the performance gain across all complexity levels. This verifies the importance of utilizing the code interface for planning instead of natural languages, which enables the model to avoid overly complex contexts and thus leads to a substantial increase in the completion rate. Additionally, the code interface also allows the debugging module to refine the planning with execution feedback, improving the efficacy of the planning process.
$\diamond$ Debugging Module. The ‘rubber duck’ debugging module enhances the performance by guiding the LLM agent to figure out the underlying reasons for the error messages. This enables EHRAgent to address the intrinsic error that occurs in the original reasoning steps. We then further illustrate the difference between debugging modules in EHRAgent and others. Self-debugging (Chen et al., 2024) that sends back the execution results with an explanation of the code for plan refinement. Reflexion (Shinn et al., 2023) sends the binary reward of whether it is successful or not back for refinement, which contains little information. In both cases, however, the error message is still information on the surface, like ‘incorrect query’, etc. This is aligned with our empirical observations that LLM agents tend to make slight modifications to the code snippets based on the error message without further debugging. Taking one step further, our debugging module in EHRAgent incorporates an error tracing procedure that enables the LLM to analyze potential causes beyond the current error message. Our debugging module aims to leverage the conversation format to think one step further about potential reasons, such as ‘incorrect column names in the query’ or ‘incorrect values in the query’.
@@
w/o medical information 36.75 28.39 6.25 30.17 47.24
w/o long-term memory 52.41 44.22 18.75 45.69 78.97
@@
w/o rubber duck debugging 50.60 46.98 12.50 47.07 70.86
@@
F.3 Cost Estimation
Using GPT-4 as the foundational LLM model, we report the average cost of EHRAgent for each query in the MIMIC-III, eICU, and TREQS datasets as \$0.60, \$0.17, and \$0.52, respectively. The cost is mainly determined by the complexity of the question (i.e., the number of tables required to answer the question) and the difficulty in locating relevant information within each table.
Appendix G Additional Empirical Analysis
G.1 Additional Question Complexity Analysis
We further analyze the model performance by considering various measures of question complexity based on the number of elements in questions, and the number of columns involved in solutions, as shown in Figure 4. Incorporating more elements requires the model to either perform calculations or utilize domain knowledge to establish connections between elements and specific columns. Similarly, involving more columns also presents a challenge for the model in accurately locating and associating the relevant columns. We notice that both EHRAgent and baselines generally exhibit lower performance on more challenging tasks [^16] Exceptions may exist when considering questions of seven elements in Figures 4(a) and 4(b), as it comprises only eight samples and may not be as representative.. Notably, our model consistently outperforms all the baseline models across all levels of difficulty. Specifically, for those questions with more than 10 columns, the completion rate of those open-loop baselines is very low (less than 20%), whereas EHRAgent can still correctly answer around 50% of queries, indicating the robustness of EHRAgent in handling complex queries with multiple elements.
G.2 Additional Error Analysis
We conducted a manual examination to analyze all incorrect cases generated by EHRAgent in MIMIC-III. Figure 6 illustrates the percentage of each type of error frequently encountered during solution generation: